OpenStudy (anonymous):

calculate the 5 unit moving average x^3-x

OpenStudy (anonymous):

ybar(x) = (1/5) sum i=-2 to 2 of (x + i h)^3 - (x + i h) = x^3 - x + 6 h^2 x